OpenStudy (loser66):

Old problem: Show that if n is odd or a multiple of 4, then 1^3 +2^3 +...+(n-1)^3 equivalent 0 (mod n) Please, help

OpenStudy (loser66):

I got stuck at the last argument.

OpenStudy (loser66):

\(\sum_{i=1}^n i^3 =\dfrac{1}{4}n^2(n+1)^2\) Hence \(\sum_{i=1}^{n-1}i^3 = \sum_{i=1}^n i^3 - n^3\) but \(n|n^3\) , hence \(\sum_{i=1}^{n-1}i^3\equiv \sum_{i=1}^n i^3\) Hence we need find the condition of n, such that \(\dfrac {1}{4}n^2 (n+1)^2\) divisible by n

OpenStudy (loser66):

We have 4 = 2^2, hence \(n\equiv 0,1\) (mod2) If \(n\equiv 0(mod2)\) , then the whole thing divisible by 4 If \(n\equiv 1(mod2)\), then \((n+1) \equiv 0(mod2)\) and the whole thing divisible by n also.

OpenStudy (loser66):

But then, I just have the condition that n is divisible by 2 only, not 4 as require. What is wrong with my argument?

OpenStudy (loser66):

@ikram002p

OpenStudy (ikram002p):

.

OpenStudy (superdavesuper):

(n(n+1)/2)^2 must be divisible by n so simplifing the expression, n(n+1)^2/4 must be an integer. That means either n/4 is an integer; i.e. n is a multiple of 4; or (n+1)^2/4 is an integer; i.e. n must be odd so that n+1 is even and (n+1)^2 will be divisible by 4.

OpenStudy (anonymous):

SPAM

OpenStudy (loser66):

Question: if n is even, we have \(\dfrac{1}{4} n*n (n+1)^4\) the first n divisible by2, same as the second one. Why do we need 4 | n?

OpenStudy (loser66):

if n is odd, n+1 is even and same argument, we have (n+1) (n+1) on the expression. Hence for all n odd, the expression divisible by 4 and that makes it be integer. Am I right?

OpenStudy (loser66):

oh, I got it.

OpenStudy (superdavesuper):

after dividing by n, the expression becomes (n*(n+1)^2)/4 and that needs to be an integer for the sum to be divisible by 4. ....not sure where the (n+1)^4 came from...

OpenStudy (superdavesuper):

oh cool :)

OpenStudy (loser66):

the first n must divisible by 4 to cancel the 4 at the bottom. And then the second one make the whole divisible by n.

OpenStudy (loser66):

for (n+1)^4, it is a mistake, it should be (n+1)^2 :) sorry.

OpenStudy (superdavesuper):

Either the first term, n, must be divisible by 4 to cancel the 4 at the bottom; OR n must be odd to make the second term divisible by 4. Then the summation is divisible by n.

OpenStudy (loser66):

Thank you so much. I do appreciate

OpenStudy (superdavesuper):

welcome n excuse me 4 not typing things out in LaTex - i dont have much luck w it on my computer